3 metal rods of same length and same area of cross section having the conductivities  1,2,3 units are connected in series. Then their effective thermal conductivity will be.....

Grade:11

2 Answers

Umakant biswal
7 years ago
@ sreenidhi 
can u remember to find out the effective resistance in the case of parallel combination of resister . 
the same formula will going to be used in this case also . 
k1 = 1 
k2 = 2 
k3 = 3
in series 
k (eq)_ = 1/k1+1/k2+1/k3 
= 1/1+1/2+1/3 = 11/6 
so, final answer will be 6/11 
beacuse 1/ k (eq ) = 11/6 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..
Talib Ahmed
5 years ago
l1+l2+l3/K=l1/k1+l2/k2+l3/k3
by using this formula,we get the final equivalent conductivity.Hope u will understand!!!